30 % of firm’s orders come from new customers. A simple random sample of 100 orders will be used to estimate the proportion of new customers. Find probability that sample proportion will be between 0.22 and 0.40
Answer by stanbon(75887)

30 % of firm’s orders come from new customers. A simple random sample of 100 orders will be used to estimate the proportion of new customers. Find probability that sample proportion will be between 0.22 and 0.40
-----------------
p = 0.3
sigma = sqrt(pq) = sqrt(0.3*0.7)
-------------------------------------
z(0.22) = (0.22-0.30)/[sqrt(0.3*0.7/100)] = -1.7457
---
z(0.40) = (0.40-0.30)/[sqrt(0.3*0.7/100)] = 2.1822
---
P(0.22 < p < 0.40) = P(-1.7457 < z < 2.1822) = 0.9450
===========================
Cheers,
Stan H.
===========================